# LoDIP: Low light phase retrieval with deep image prior

### v1

@@ Abstract
Phase retrieval (PR) is a fundamental challenge in scientific imaging, enabling nanoscale techniques like coherent diffractive imaging (CDI). Imaging at low radiation doses becomes important in applications where samples are susceptible to radiation damage. However, most PR methods struggle in low dose scenario due to the presence of very high shot noise. Advancements in the optical data acquisition setup, exemplified by in-situ CDI, have shown potential for low-dose imaging. But these depend on a time series of measurements, rendering them unsuitable for single-image applications. Similarly, on the computational front, data-driven phase retrieval techniques are not readily adaptable to the single-image context. Deep learning based single-image methods, such as deep image prior, have been effective for various imaging tasks but have exhibited limited success when applied to PR. In this work, we propose LoDIP which combines the in-situ CDI setup with the power of implicit neural priors to tackle the problem of single-image low-dose phase retrieval. Quantitative evaluations demonstrate the superior performance of LoDIP on this task as well as applicability to real experimental scenarios.
@@
deep image prior, deep generative models, computational imaging, phase retrieval, inverse problems, low light imaging, low photon count
@@ I Introduction
Figure 1: Coherent Diffraction Imaging (CDI) employs a coherent X-ray beam directed at a sample, capturing the resulting diffraction pattern on a 2D detector. A computational algorithm is then applied to reconstruct the desired sample image. Inspired by in-situ CDI [1], LoDIP introduces two modifications to the CDI setup. First, it involves imaging the sample alongside a static region. Secondly, the static region is exposed to a high radiation dose, while the sample’s exposure is significantly reduced through the use of an attenuator. In the computational step, LoDIP takes both the diffraction pattern and an estimated reconstruction of the static region as inputs, generating a sample reconstruction as its output.
Coherent diffractive imaging (CDI) is a lensless imaging technique [2] used for high resolution imaging at nanoscale. CDI has found broad applications across different disciplines due to its remarkable ability to provide high-resolution structural information about a wide range of specimens from biological specimens to nanoscale objects [3].
@@
Unlike visible light, X-rays have high penetrating power and thus can be used to image thick, unfixed specimens. However, many samples of interest for CDI, such as biological material, polymers or organic semiconductors, require minimal radiation exposure to prevent damage during data acquisition [4, 5]. Thus, there is considerable interest in techniques that minimize sample radiation exposure and enable X-ray imaging at extremely low photon counts.
The main challenge is that imaging at low photon counts leads to very high shot noise in the acquired measurements (diffraction pattern) and makes the subsequent image reconstruction (i.e. phase retrieval) very challenging. Under high shot noise, iterative phase retrieval algorithms, including ER, HIO, become unstable, facing challenges such as getting trapped in local minima, stagnation, and failure to converge.
@@
One category of proposed solutions for the low-dose challenge involve modifications to the optical data acquisition setup of CDI [6, 7, 1]. A notable example is in-situ CDI, as introduced by Lo et al. [1] and further explored in [8]. A dose-tolerant static region is positioned next to the sample and is illuminated with a higher radiation dose. This ensures sufficient light reaches the detector while maintaining a low radiation dose on the sample. Multiple measurements are taken over time and the static region is leveraged as a robust time-overlap constraint which regularizes the phase retrieval optimization. However, in order to leverage the time-invariant static region as real-space constraint, this method requires multiple measurements (that is multiple diffraction patterns). Adapting this setup for single-image scenarios results in suboptimal reconstructions (see HIO-stat column in Fig. 4 ).
The solution we propose in this paper is to design a novel data-driven computational algorithm to address single image low-dose phase retrieval. Data-driven approaches have shown potential for simpler versions of phase retrieval at low-photon counts [9, 10], but they cannot be readily adapted to the single-image context due to the requirement for a large training dataset. Deep learning based single-image methods such as deep image prior (DIP) [11, 12] have been used in the past in the single-image context. While they have demonstrated success in simpler phase retrieval scenarios [13, 14, 15], they struggle in low dose situations, as evident from our experiments Fig. 4 and from previous literature on the topic [16, 17].
To address these issues, in this work we propose a deep learning method designed for single-image low-dose phase retrieval. More specifically, we propose LoDIP, a single-image method for low-dose phase retrieval based on DIP and inspired by in-situ CDI. We use the modified imaging setup consisting of a static region illuminated with a high dose of radiation alongside the sample of interest which is illuminated by low-light. This increases the amount of light incident on the detector, without increasing the radiation dose on the sample. Then we modify the DIP framework to exploit the additional constraints coming from this setup. Experiments show that the proposed method LoDIP can reliably work for single-image low-dose phase retrieval both for simulated and experimental data. In particular our experiments show the advantages of LoDIP over state of the art methods: first, being completely data driven, LoDIP does not require specialized algorithmic training to optimize the parameters to obtain satisfactory results. This makes LoDIP accessible to a broad user community. Secondly, our experiments demonstrate that LoDIP outperforms state-of-the-art methods by yielding reconstructions with reduced noise (higher peak signal-to-noise ration) while preserving resolution (comparable Fourier Ring Correlation). Third, since LoDIP is based on the DIP framework, it does not require a large training dataset, in contrast with other deep learning approaches. Finally, being designed for single-image phase retrieval, LoDIP can produce comparable reconstructions to the original in situ CDI method without requiring a time series of measurements.
@@ II-A Coherent Diffractive Imaging
To meet the oversampling criteria, which helps to mitigate the non-uniqueness of the phase retrieval problem, the original object $\mathbf{X}$ is placed in an empty background such that $m=2\times n$. In simulated settings, this oversampling is achieved by zero-padding $\mathbf{X}$ before applying the forward operator. The overall goal of the phase retrieval problem is to recover the image $\mathbf{X}$ from the captured diffraction pattern $\mathbf{Y}$.
Generally, some information about the support of the object (i.e. the location of the object) within the empty background is known. Let $\mathbf{S}_{0}\in\mathbb{C}^{m\times m}$ be the known support information. $S_{0}$ is a matrix containing ones where the object of interest is estimated to be, zeros otherwise. In general one may not know the exact support of the object of interest, but only an estimate of it. This makes the retrieval harder since it leaves some translation freedom (see Fig. 6 for experiments with inaccurate support estimate).
@@
where the first term imposes that II.1 is satisfied, while the second term is the support constraint. The objective function in the above formulation is referred to as the magnitude constraint or data consistency term. The support constraint imposes that the reconstructed object has the estimated support. We note that the objective function can also be written in terms of the Fourier magnitudes (not squared), i.e. $l(\sqrt{\mathbf{Y}},|\mathcal{F}(\mathbf{X})|)$, this gave similar results in our experiments as the squared case.
Figure 2: Examples from the simulated data at different levels of illumination of the sample region as measured in photons per pixel (Np/px). [Left] The first two columns are the diffraction pattern and the image sample without a static structure. [Right] And the last two columns are the same sample with a static region. [Top] The first row has a high illumination of 80k Np/px on both the sample and the static region. [Bottom] Whereas the second row has a high illumination of 80k Np/px on the static region and a low illumination of 800 Np/px on the sample. It can be seen that at lower illumination (bottom row), the sample has very low pixel values.
@@ II-B Phase retrieval for low-light imaging
Imaging at cellular or atomic scales necessitates use of X-ray radiation. However, many samples of interest for CDI, such as biological material, polymers or organic semiconductors, require minimal radiation exposure to prevent damage during data acquisition. As a consequence,there is considerable interest in techniques that minimize sample radiation exposure and enable X-ray imaging at extremely low-photon counts. The major challenge of working in the low-dose case is the strong presence of noise in the captured signal. In practice, most of the existing PR methods [20, 21, 22, 23, 24, 25] struggle to work in the low dose setting (high noise). This comes in addition to the existing challenge of working without accurate support information which itself is sufficient for many of the existing algorithms to fail (unless additional strategies are adopted like the shrink-wrap method [26]).
@@
Attempts have been made at low-photon phase retrieval that modify the imaging setup of the problem [6, 7, 1]. While [6] relies on a phase diverse approach, [7, 1] use in the imaging setup a static region made of heavy metals (usually gold) which can withstand high energy light. Among these methods, in-situ CDI [1] achieves a substantial order of magnitude dose reduction over other methods. The main idea is to monitor a dynamic process over time containing two distinct regions: a dynamic area containing the sample which continuously changes, and a static region that remains stationary in time. The static region is subject to significantly higher illumination compared to the sample, ensuring sufficient light falls on the detector, while keeping the radiation dose low for the sample. A series of diffraction patterns is collected over time, resulting in interference between the static and dynamic regions. The static region’s presence in each time step’s measurement offers a strong overlap-in-time constraint for facilitating the convergence of phase retrieval optimization. In the original in situ CDI paper experiments show the reconstruction of the motion of low-dose samples such as glioblastoma cells through Oversampling Smoothness (OSS) algorithm [24].
II-C Deep Learning for phase retrieval
Recent data-driven deep learning methods for phase retrieval [27, 28, 10, 29] have demonstrated that deep neural networks (DNN) within a supervised learning framework can enhance performance in low-light conditions. However, despite some initial successes, deep learning methods for phase retrieval are yet to see widespread adoption among microscopy practitioners. There are a few reasons for why this happens: model training often requires a large dataset, making them unusable in the single-image setting; training often a long time and GPU resources; the reconstruction performance heavily depends on the quality of data provided for the model training. Finally, each model has to be retrained to be used on samples captured under different experimental settings. In this work, we propose a deep learning-based method based on deep image prior which can function in a single-image setting without requiring a large dataset and can be used across datasets and experimental setups.
II-D Neural Networks as a prior
After the success of convolutional neural networks (CNNs) on image classification tasks [30], neural networks and particularly CNNs have improved the state-of-the-art in many computer vision tasks. The majority of these instances were within the framework of supervised learning, requiring the use of extensive datasets for training. More recent works (see for instance [31, 32, 33]) showed that neural networks can succeed in approximating images distribution (this is known as generative modeling). This success led to the use of trained CNNs as a prior for regularizing inverse problems in computer vision [34]. The main idea is to pre-train the model on a similar dataset (possibly for a different task) and use them for inverse problems in a single-sample setting, i.e. without re-training. Around the same time it was discovered that even an untrained neural network can be used as a strong prior for natural images [11, 12] and can be successfully used to regularize inverse problems on natural images.
The surprising success of untrained neural prior or deep image prior (DIP) on inverse problems in image restoration and enhancement has been followed by works aimed at understanding its underlying mechanisms [35]. Subsequent research has resulted in various extensions and adaptations, including approaches involving early stopping strategies [36, 14], methods addressing spectral bias improvements [37], and innovative applications [38]. For a comprehensive overview of the extensive body of work on untrained neural priors in image enhancement, we recommend a recent comprehensive survey [39].
Untrained neural priors have also been applied to computational imaging inverse problems [40, 41, 42] including simpler versions of phase retrieval [13, 14, 15, 43]. However, compared to the setting used in this paper (far-field Fraunhoffer diffraction), these works tackle a much simpler settings for phase retrieval (e.g. near-field Fresnel diffraction).
In this work, we adopt the in-situ CDI setup for the single-image phase retrieval and complement it with an untrained neural network prior in the computational algorithm. Our method is able to produce accurate reconstructions even in the low dose setting, thanks to the in-situ CDI set up and is able to produce results for single-image phase retrieval since the untrained neural network does not require any data. Experiments on simulated and experimental data demonstrate the efficacy of this method over either of these methods alone and other state of the art methods.
@@ III Proposed method
The proposed method, which we refer to as LoDIP, modifies the data acquisition setup of conventional CDI to incorporate a high-dose static region inspired by the in-situ CDI setup. While in-situ CDI leverages the static region as a time-invariant constraint to reconstruct a sequence of dynamic measurements, the high-dose static region serves a different purpose in LoDIP where we are interested in the single-image setting.
@@
While these adaptations significantly enhance the performance of established phase retrieval methods (such as HIO) at low radiation doses, our experiments (see Fig. 3), show that LoDIP can further improve the final reconstruction quality by incorporating an untrained neural prior for phase retrieval.
III-A LoDIP: Data acquisition
The sample of interest is placed within a finite support next to a static region of heavily scattering, dose-tolerant object such as a gold (Au) pattern on an optical stage. The X-ray illumination on the dose-sensitive sample is reduced to a tolerable limit by the presence of an attenuator, while the static region is exposed to the full dose of the incident illumination. Far-field diffraction patterns recorded from this setup are formed by the interference in Fourier space between the high-dose static region and the sample. These contain Poisson noise relative to the total illumination on the detector. Given the high-dose illumination on the static region and the known support, it is possible to obtain a high-quality static region reconstruction. In this work we use an established iterative method, Generalized Proximal Smoothing (GPS) initialized with 1000 iterations of HIO.
We note explicitly that the placement of the object and the static structure do not influence the reconstruction given by LoDIP, as long as their supports do not overlap. More specifically, let $\mathcal{S}\left(\cdot\right)$ be the operator which extracts the support (location of non-zero elements) of a given image. Then, we assume that the object and the static region have a non-overlapping supports, i.e. $\mathcal{S}\left(\mathbf{X}\right)\odot\mathcal{S}\left(\mathbf{U}\right)=[\mathbf{0}]_{m\times m}$.

### v2

# Low-light phase retrieval with implicit generative priors

@@ Abstract
Phase retrieval (PR) is fundamentally important in scientific imaging and is crucial for nanoscale techniques like coherent diffractive imaging (CDI). Low radiation dose imaging is essential for applications involving radiation-sensitive samples. However, most PR methods struggle in low-dose scenarios due to high shot noise. Recent advancements in optical data acquisition setups, such as in-situ CDI, have shown promise for low-dose imaging, but they rely on a time series of measurements, making them unsuitable for single-image applications. Similarly, data-driven phase retrieval techniques are not easily adaptable to data-scarce situations. Zero-shot deep learning methods based on pre-trained and implicit generative priors have been effective in various imaging tasks but have shown limited success in PR. In this work, we propose low-dose deep image prior (LoDIP), which combines in-situ CDI with the power of implicit generative priors to address single-image low-dose phase retrieval. Quantitative evaluations demonstrate LoDIP’s superior performance in this task and its applicability to real experimental scenarios.
@@
deep generative models, computational imaging, phase retrieval, inverse problems, low light imaging, low photon count, zero-shot learning
@@ I Introduction
Coherent diffractive imaging (CDI) is a lensless imaging technique [1] used for high resolution imaging at nanoscale. CDI has found broad applications across different disciplines due to its remarkable ability to provide high-resolution structural information about a wide range of specimens from biological to nanoscale objects [2]. Unlike visible light, X-rays have high penetrating power and thus can be used to image thick, unfixed specimens. However, many samples of interest for CDI, such as biological materials, polymers or organic semiconductors, require minimal radiation exposure to prevent damage during data acquisition [3, 4]. Thus, there is considerable interest in techniques that minimize sample radiation exposure and enable X-ray imaging at extremely low photon counts. However, imaging at low photon counts is very challenging as it leads to very high shot noise in the acquired measurements (diffraction pattern) and makes the subsequent image reconstruction (i.e. phase retrieval) very hard. Under high shot noise, iterative phase retrieval algorithms, including ER, HIO, become unstable, as they often get trapped in local minima, stagnate, and fail to converge [5, 6].
@@
Motivation. One category of proposed solutions for the low-dose challenge involve modifications to the optical data acquisition setup of CDI [7, 8, 9]. A notable example is in-situ CDI, as introduced by Lo et al. [9] and further explored in [10]. A dose-tolerant static region is positioned next to the sample and is illuminated with a higher radiation dose. This ensures sufficient light reaches the detector while maintaining a low radiation dose on the sample. Multiple measurements are taken over time and the static region is leveraged as a robust time-overlap constraint which regularizes the phase retrieval optimization. However, in order to leverage the time-invariant static region as real-space constraint, this method requires multiple measurements (that is multiple diffraction patterns). Adapting this setup for single-image scenarios results in suboptimal reconstructions (see HIO-stat column in Fig. 4 ).
On the computational front, data-driven supervised learning approaches have shown potential for simpler versions of phase retrieval at low-photon counts [11, 12, 13]. However, the necessity for a substantial dataset is a challenge to meet for numerous scientific imaging applications. Furthermore, these methods are prone to generalization problems if the training data diverges from the target domain. Consequently, deep learning-based zero-shot methods such as deep image prior (DIP) [14, 15] have been employed. These do not require pre-training, rather the primary concept is to leverage the inductive bias induced by the neural network architecture as a prior for natural images. We refer to this class of methods as implicit generative priors to highlight the absence of training data. While successful in simpler phase retrieval scenarios [16, 17, 18], these methods often struggle to converge to satisfactory solutions in more challenging realistic scenarios, such as the single-shot far-field Phase Retrieval (FFPR) studied in this work. This observation is supported by our low-dose experiments (see Fig. 4) and corroborated by prior literature [19, 20]. The difficulty can be attributed to the well-known trivial ambiguities in the phase retrieval problem [5].
Basic Idea. LoDIP, a single-image method for low-dose phase retrieval, modifies both the experimental setup and computational algorithm of classical CDI. We modify the imaging setup of CDI to incorporate a static region illuminated with a high dose of radiation alongside the sample of interest which is illuminated by a low dose. Then we modify the DIP framework to exploit the additional physical constraints coming from this setup. The high-dose static region increases the amount of light incident on the detector, without increasing the radiation dose on the sample. This results in a reduction of noise in the captured diffraction pattern. Moreover, the static region aids in the convergence of implicit generative prior by alleviating the trivial ambiguities [21].
Innovation. LoDIP introduces innovation by concurrently modifying both experimental setup and computational algorithm. By integrating physics-based constraints from the modified experimental setup, LoDIP effectively resolves instability and artifacts typical of conventional DIP methods in single-shot far-field phase retrieval. This ensures stable optimization and precise reconstruction from low-light diffraction patterns, leveraging the benefits of deep image prior optimization while overcoming inherent challenges in single-shot far-field phase retrieval.
Unlike in-situ CDI setup, which requires multiple diffraction patterns, LoDIP operates in a single-image framework. Additionally, in contrast to in-situ CDI, where the static region is used as a time-invariant constraint present in multiple measurements, in LoDIP, the high-dose imaging of the static region serves two novel purposes: (a) being imaged at high-dose, it provides an accurate reconstruction of the static region, which acts as a strong constraint crucial for achieving a high-quality reconstruction of the sample; and (b) enhancing the convergence of phase retrieval optimization by mitigating ambiguities arising from symmetries in the forward process, a significant challenge in phase retrieval [5, 21].
Contributions. Experiments demonstrate that LoDIP is effective for both simulated and experimental data. The LoDIP framework is very flexible, compatible with any implicit generative prior method, and can be tailored to diverse experimental setups. Unlike iterative phase retrieval algorithms, LoDIP does not necessitate specialized tuning to optimize the algorithmic parameters for satisfactory results, making LoDIP a user-friendly tool for a wide user community. The contributions of this work can be summarized as follows:
  • A zero-shot method for low-dose phase retrieval. LoDIP uses the inductive bias of an implicit generative prior and does not require large dataset for training.
  • Extensive experiments on simulated and experimental datasets demonstrate that LoDIP outperforms state-of-the-art methods across a range of image quality metrics.
  • On the experimental side, LoDIP being designed for single-image phase retrieval, can produce reconstructions comparable to the original in situ CDI method without requiring a time series of measurements.
@@ II-A Coherent Diffractive Imaging
Generally, some information about the support of the object (i.e. the location of the object) within the empty background is known. Let $\mathbf{S}_{0}\in\mathbb{C}^{m\times m}$ be the known support information. $S_{0}$ is a matrix containing ones where the object of interest is estimated to be, zeros otherwise. In general one may not know the exact support of the object of interest, but only an estimate.
@@
where the first term imposes that II.1 is satisfied (magnitude constraint), while the second term is the support constraint.
Figure 1: Examples from the simulated data at different levels of illumination of the sample region as measured in photons per pixel (Np/px). [Left] The first two columns are the diffraction pattern and the image sample without a static structure. [Right] And the last two columns are the same sample with a static region. [Top] The first row has a high illumination of 80k Np/px on both the sample and the static region. [Bottom] Whereas the second row has a high illumination of 80k Np/px on the static region and a low illumination of 800 Np/px on the sample. At lower illumination levels (bottom row), the sample exhibits significantly reduced pixel values. Lesser illumination results in higher presence of Poisson noise. Thus different illumination levels correspond to different levels of noise.
@@ II-B Phase retrieval for low-light imaging
Challenge The major challenge of imaging at low-photon counts is the strong presence of noise in the captured signal. In practice, most of the existing PR methods [23, 24, 25, 26, 27, 6] struggle to work in the low dose (high noise) setting. This comes in addition to the existing challenge of working without accurate support information which itself is sufficient for many of the existing algorithms to fail.
@@
Existing solutions Attempts have been made at low-photon phase retrieval that modify the imaging setup of the problem [7, 8, 9]. While [7] relies on a phase diverse approach, [8, 9] use in the imaging setup a static region made of heavy metals (usually gold) which can withstand high energy light. Among these methods, in-situ CDI [9] achieves a substantial order of magnitude dose reduction over other methods. However, in-situ CDI uses a time-series of diffraction patterns which should contain a time-invariant ”static” region. The static region’s presence in each time step’s measurement offers a strong overlap-in-time constraint for facilitating the convergence of phase retrieval optimization. Fourier holography [28, 29] also utilizes two regions of interest. However, unlike Fourier holography, LoDIP can handle varying illuminations and doesn’t impose geometric constraints between the regions.
II-C Data-driven phase retrieval
Data-driven methods have shown promise for phase retrieval [30, 12, 31] including at low-dose [32]. Subsequently, the end-to-end supervised learning paradigm has been extended by incorporating known physical models in the neural-network training. This includes utilizing an unrolled iterative algorithm [33, 34] or by devising a self-supervised loss function [35, 36] capable of learning from measurements alone or using initial phase estimates from a physical model as inputs to the network [37, 38]. For a comprehensive review of data-driven phase retrieval, we refer the reader to [39].
However, these methods are not yet widely adopted by microscopy practitioners due to several challenges: they often require large datasets for training, extensive computational resources, and their performance is heavily reliant on the quality of the training data. Additionally, retraining is necessary to adapt the model to different experimental settings. In this study, we propose a single-image deep learning method based on deep image prior that overcomes these limitations, requiring no large dataset and being adaptable across various experimental setups.
II-D Zero-shot deep learning for phase retrieval
Pre-trained generative priors. Following the breakthroughs achieved by deep generative models in approximating image distributions [40, 41, 42, 43], such models have been leveraged as priors to regularize visual inverse problems [44, 43], including phase retrieval [13]. Methods based on pre-trained generative priors are commonly referred to as “zero-shot” as no additional task-specific training is performed in these methods. Despite being unsupervised and not reliant on labeled datasets, these approaches still necessitate pretraining on extensive datasets. Additionally, akin to supervised learning methods, these pre-trained generative priors face challenges in generalization when the training data deviates from the target domain.
Implicit generative priors. A parallel line of work discovered that even a convolutional neural network (CNN) without training can serve as a robust prior for natural images, and can be effectively employed to regularize inverse problems on such images. Beginning with the so-called deep image prior (DIP) [14]. This has paved the way for a suite of methods including implicit neural representations (INR) [45, 15], deep decoder [17], double-DIP [46], random projection based methods [47] and enhanced versions of DIP [48]. For an exhaustive review, readers are directed to [49]. In the absence of an overarching terminology for these methods, we label them as ‘implicit generative priors’. This term highlights the inductive bias derived from the network architecture rather than the training data. It also differentiates these methods from the use of pre-trained generative priors. Both types are frequently referred to as ’zero-shot’ deep learning methods. Finally, it highlights the generative nature of our algorithm which takes noise as input and generates a natural image.
For Phase Retrieval. Several of these have also been applied to computational imaging inverse problems [50, 51, 52] including simpler versions of phase retrieval [16, 17, 18, 53].  [^1] Certain self-supervised data-driven methods have demonstrated efficacy even with a single sample [35, 54], resembling a zero-shot approach akin to an implicit generative prior. Instead of a random input seed, these methods rely on a sample-dependent input [39]. However, the settings addressed in these works are considerably simpler compared to the single-shot far-field (Fraunhoffer) diffraction with high noise scenario considered in this paper. A naive application of DIP and related methods is known to struggle in such realistic phase retrieval scenarios [53, 20] .
@@ III Proposed method
Figure 2: Coherent Diffraction Imaging (CDI) employs a coherent X-ray beam directed at a sample, capturing the resulting diffraction pattern on a 2D detector. A computational algorithm is then applied to reconstruct the desired sample image. (Left) Inspired by in-situ CDI [9], LoDIP introduces two modifications to the CDI setup. First, it involves imaging the sample alongside a static region. Secondly, the static region is exposed to a high radiation dose, while the sample’s exposure is significantly reduced. A customized sample grid and holder is used with an attenuator placed on top of the sample to reduce the incident radiation dose on the sample. (Center) In the computational step, LoDIP takes both the diffraction pattern and an estimated reconstruction of the static region as inputs, generating a sample reconstruction as its output. (Right) LoDIP uses the output of an implicit generative model $g_{w}$ as an estimate of the sample $\hat{X}$ and iteratively updates the generator parameters $w$ to refine the estimate.
The proposed method, which we refer to as LoDIP, modifies the data acquisition setup of conventional CDI to incorporate a high-dose static region inspired by the in-situ CDI setup. While in-situ CDI leverages the static region as a time-invariant constraint to reconstruct a sequence of measurements capturing a dynamic process, the high-dose static region serves a different purpose in LoDIP where we are interested in the single-image setting.
@@
While these adaptations significantly enhance the performance of established phase retrieval methods (such as HIO) at low radiation doses, our experiments (see Fig. 3), show that LoDIP can further improve the final reconstruction quality by incorporating an implicit generative prior for phase retrieval.
III-A LoDIP: experimental setup
The sample of interest is placed within a finite support next to a static region of heavily scattering, dose-tolerant object such as a gold (Au) pattern on an optical stage. The X-ray illumination on the dose-sensitive sample is reduced to a tolerable limit by the presence of an attenuator mounted to the sample support (see Figure 2), while the static region is exposed to the full dose of the incident illumination. The design of the customized sample support would need to include a SiN attenuator over the dose-sensitive region with a thickness tailored to the desired dose reduction factor and the energy of the incident x-ray beam. For example, to achieve a dose reduction factor of three orders of magnitude at a soft x-ray beamline operating within the water window as considered in our simulations ($\sim 525eV$), a $Si_{3}N_{4}$ thickness of roughly 1.5 $\mu$m is required  [^2] This estimated value was obtained using the tabulated atomic scattering factors for solid $Si_{3}N_{4}$ [56, 57] .
Far-field diffraction patterns recorded from this setup are formed by the interference in Fourier space between the high-dose static region and the sample. These contain Poisson noise relative to the total illumination on the detector. Given the high-dose illumination on the static region and the known support, it is possible to obtain a high-quality static region reconstruction. In this work we use an established iterative method, Generalized Proximal Smoothing (GPS) initialized with 1000 iterations of HIO.
The placement of the object and the static structure do not influence the reconstruction given by LoDIP, as long as their supports do not overlap, i.e. $\mathcal{S}\left(\mathbf{X}\right)\odot\mathcal{S}\left(\mathbf{U}\right)=[\mathbf{0}]_{m\times m}$, where $\mathcal{S}\left(\cdot\right)$ extracts the support the image.